# Searches for the BSM scenarios at the LHC using decision tree based machine learning algorithms: A comparative study and review of Random Forest, Adaboost, XGboost and LightGBM frameworks

### v1

@@ 5.1 Cut-and-count analysis
From the cut optimization, we have found that distinct sets of cuts become necessary for various benchmark points due to their varying $\Delta m$ values. For benchmark points with substantial $\Delta m$, cuts involving solely $p_{T}^{l_{1,2,3}}$, $m_{SFOS}$, $m_{T}$, $\rm E{\!\!\!/}_{T}$, number of jets ($N_{j}$) and $\gamma$ variables prove adequate. Here, the same cut set is effective for both the benchmark points with $\Delta m=700,1100$. Three different signal regions are defined in the Table 4 as SR-A, SR-B and SR-C corresponding to $\Delta m=50$ GeV, $\Delta m=70$ GeV and $\Delta m=700$ 1100 GeV. We provide the signal yield, individual background yields, and total background yield, along with the signal significance without uncertainty ($\sigma_{ss}^{0}$) and with 10% ($\sigma_{ss}^{10}$) and 20% uncertainty ($\sigma_{ss}^{20}$) in Table 5. The signal significance without and with uncertainty are calculated using the formulas mentioned in the Equations. 9 and 10 (see section. 2). The $\sigma_{ss}^{0}$ ($\sigma_{ss}^{20}$) values corresponding to BP1, BP2, BP3, and BP4 are 4.72 (0.19), 2.04 (0.07), 10.77 (7.98), and 1.59 (1.30), respectively. The cut-based method shows that BP1, BP2 and BP3 are within reach of HL-LHC [^13] Due to the presence of huge background or smaller $S/B$ ratio, the compressed scenario (BP1, BP2) will not be sensitive even we consider a 10% systematic uncertainty..
@@ 5.2.2 Feature importance with SHapley
As discussed in Sec.5.2, for this analysis, more than 20 distinct features have been used to train the ML model and the impact of these features in discriminating the signal and background can be determined using SHAP package Shapley+1953+307+318 ; lundberg2018consistent ; pedregosa2018scikitlearn . In the context of tree ensemble methods like gradient boosting methods or random forests, it is common to assign importance values to each input feature in order to comprehend the predictions. These values can be calculated for an individualized prediction or the global prediction of an entire dataset. Popular packages like XGBoost Shapley+1953+307+318 ; lundberg2018consistent ; pedregosa2018scikitlearn offer implementations of tree ensembles that enable users to calculate a metric of feature importance. These metrics aim to condense the complexity of ensemble models and offer an understanding of the key features influencing the model’s predictions. Global feature importance is derived through three main methodologies gain, split count, permutation and Saabas method. However, most of these methods of calculating feature importance metrics are inconsistent and ineffective.
@@ 6 Summary
Over the years, machine laerning techniques have significantly boosted efficiency and accuracy in HEP data analysis for both experimental and phenomenological works. For the analyses of event triggering, jet tagging, particle identification, event selection, object reconstruction, event classification etc., the HEP community has been widely using the boosted decision tree (BDT) algorithms for a long time. We have briefly summarized the major important analyses performed by the ATLAS and CMS collaboration utilizing LHC data as well as the works of the phenomenological groups involving the use of ML algorithms with emphasis on DT-based approaches. We have also outlined the basic concepts of machine learning along with different kinds of loss functions and their roles, issues of underfitting, overfitting, etc. Evaluation of the effectiveness and accuracy of the ML models in solving a specific task, e.g., classifying signal and background events, can be done using performance metrics. Different kinds of metrics relevant to particle physics analysis, like ROC curve, F-score, ams score, etc., are discussed in detail.
@@
In this article, we have focused solely on a prominent machine learning technique, namely Decision Tree (DT), specifically boosted decision tree, which is primarily used for tasks in supervised classification and regression. We particularly explore four decision tree-based ML algorithms, namely, Random Forest, AdaBoost and two gradient boosting frameworks such as XGBoost, and LightGBM, in the context of Supersymmetry which is a well-promising candidate for beyond the Standard model framework. We summarize the basic concepts and working principles of these four algorithms along with the flowcharts. In numerous analyses, both the ATLAS and CMS collaborations, HEP phenomenological groups have employed BDT algorithms to enhance the sensitivity of sparticle searches within the framework of RPC and RPV SUSY models. However, exploring the compressed SUSY parameter space still reamsins as a significant challenge.
@@
Using an example of wino type electroweakino productions at the high luminosity LHC, we demonstrate how these algorithms lead to improvement in the search sensitivity compared to traditional cut-based methods in both compressed and non-compressed R-Parity conserving SUSY scenarios. The optimization of hyperparameters and its role in signal significance are studied in detail for these four algorithms. We have also discussed how to find out the individualized and global feature importance in ML methods using SHAP package. We have found that there are $\sim$ 30-60% gains in the significance for the signal benchmark points with a large mass gap between NLSP-LSP pair, whereas, for the compressed region (mass gap 50 and 70 GeV), the signal significance improves by $\sim$ 3-4 times compared to the cut-based analysis. We have obtained that the LightGBM and XGBoost algorithms perform better than the other two algorithms Random Forest and AdaBoost. Although the deep neural networks (DNN) or deep learning (DL) techniques, which are based on multilayer NN, are gaining popularity, BDTs continue to be highly relevant and important in High Energy Physics due to their user-friendly nature, interpretability, computational efficiency, and more.

### v2

@@ 5.1 Cut-and-count analysis
From the cut optimization, we have found that distinct sets of cuts become necessary for various benchmark points due to their varying $\Delta m$ values. For benchmark points with substantial $\Delta m$, cuts involving solely $p_{T}^{l_{1,2,3}}$, $m_{SFOS}$, $m_{T}$, $\rm E{\!\!\!/}_{T}$, number of jets ($N_{j}$) and $\gamma$ variables prove adequate. Here, the same cut set is effective for both the benchmark points with $\Delta m=700,1100$. Three different signal regions are defined in the Table 4 as SR-A, SR-B and SR-C corresponding to $\Delta m=50$ GeV, $\Delta m=70$ GeV and $\Delta m=700$ 1100 GeV. We provide the signal yield, individual background yields, and total background yield, along with the signal significance without uncertainty ($\sigma_{ss}^{0}$) and with 10% ($\sigma_{ss}^{10}$) and 20% uncertainty ($\sigma_{ss}^{20}$) in Table 5. The signal significance without and with uncertainty are calculated using the formulas mentioned in the Equations. 10 and 11 (see section. 2). The $\sigma_{ss}^{0}$ ($\sigma_{ss}^{20}$) values corresponding to BP1, BP2, BP3, and BP4 are 4.72 (0.19), 2.04 (0.07), 10.77 (7.98), and 1.59 (1.30), respectively. The cut-based method shows that BP1, BP2 and BP3 are within reach of HL-LHC [^12] Due to the presence of huge background or smaller $S/B$ ratio, the compressed scenario (BP1, BP2) will not be sensitive even we consider a 10% systematic uncertainty..
@@ 5.2.2 Feature importance with SHapley
As discussed in Sec.5.2, for this analysis, more than 20 distinct features have been used to train the ML model and the impact of these features in discriminating the signal and background can be determined using SHAP package Shapley+1953+307+318 ; lundberg2018consistent ; pedregosa2018scikitlearn . In the context of tree ensemble methods like gradient boosting methods or random forests, it is common to assign importance values to each input feature in order to comprehend the predictions. These values can be calculated for an individualized prediction or the global prediction of an entire dataset. Popular packages like XGBoost Shapley+1953+307+318 ; lundberg2018consistent ; pedregosa2018scikitlearn offer implementations of tree ensembles that enable users to calculate a metric of feature importance. These metrics aim to condense the complexity of ensemble models and offer an understanding of the key features influencing the model’s predictions. Global feature importance is derived through three main methodologies: gain, split count, permutation, and Saabas method. However, most of these methods of calculating feature importance metrics are inconsistent and ineffective.
@@ 6 Summary
Over the years, machine learning techniques have significantly boosted efficiency and accuracy in HEP data analysis for both experimental and phenomenological works. For the analyses of event triggering, jet tagging, particle identification, event selection, object reconstruction, event classification etc., the HEP community has been widely using the boosted decision tree (BDT) algorithms for a long time. We have briefly summarized the major important analyses performed by the ATLAS and CMS collaboration utilizing LHC data, as well as the works of the phenomenological groups involving the use of ML algorithms with emphasis on DT-based approaches. We have also outlined the basic concepts of machine learning along with different kinds of loss functions and their roles, issues of underfitting, overfitting, etc. Evaluation of the effectiveness and accuracy of the ML models in solving a specific task, e.g., classifying signal and background events, can be done using performance metrics. Different kinds of metrics relevant to particle physics analysis, like ROC curve, F-score, ams score, etc., are discussed in detail.
@@
In this article, we have focused solely on a prominent machine-learning technique, namely Decision Tree (DT), specifically boosted decision tree, which is primarily used for tasks in supervised classification and regression. We particularly explore four decision tree-based ML algorithms, namely, Random Forest, AdaBoost and two gradient boosting frameworks such as XGBoost, and LightGBM, in the context of Supersymmetry which is a well-promising candidate for beyond the Standard model framework. We summarize the basic concepts and working principles of these four algorithms along with the flowcharts. In numerous analyses, both the ATLAS and CMS collaborations, HEP phenomenological groups have employed BDT algorithms to enhance the sensitivity of sparticle searches within the framework of RPC and RPV SUSY models. However, exploring the compressed SUSY parameter space still remains as a significant challenge.
@@
Using an example of wino-type electroweakino productions at the high luminosity LHC, we demonstrate how these algorithms lead to improvement in the search sensitivity compared to traditional cut-based methods in both compressed and non-compressed R-Parity conserving SUSY scenarios. The optimization of hyperparameters and its role in signal significance are studied in detail for these four algorithms. We have also discussed how to find out the individualized and global feature importance in ML methods using SHAP package. We have found that there are $\sim$ 30-60% gains in the significance for the signal benchmark points with a large mass gap between NLSP-LSP pair, whereas, for the compressed region (mass gap 50 and 70 GeV), the signal significance improves by $\sim$ 3-4 times compared to the cut-based analysis. We have obtained that the LightGBM and XGBoost algorithms perform better than the other two algorithms Random Forest and AdaBoost. Although the deep neural networks (DNN) or deep learning (DL) techniques, which are based on multilayer NN, are gaining popularity, BDTs continue to be highly relevant and important in High Energy Physics due to their user-friendly nature, interpretability, computational efficiency, and more.